OpenStudy (andriani):

as a food preservative, citric acid is mixed with sodium citrate which is used as a buffer solution. calculate how much ph of foods containing citric acid 0.1M and sodium nitrate 0.3M (Ka citric acid 8x10 ^ -4)

OpenStudy (gebooors):

Use HEnderson-Hasselbalch equation pKa = - log (Ka)

OpenStudy (andriani):

please show me the details

OpenStudy (zale101):

Henderson-Hasselbalch equation states \(pH=pK_a+log[\frac{[A^-]}{[HA]}]\) Giving the balanced euqation \( HA\rightarrow\ H^+(aq)+A^-(aq)\) HA is the generic acid. H+ is the hydrogen ion that lost an e- and A- is the conjugate base of the acid HA [A-] is the concentration of the conjugated acid [HA] is the concentration of the acid To get the pKa use \(pKa=-log(Ka)\) Good thing they gave you the Ka. All you have to do now is take the negative log of it and you will get the pKa. Follow this balanced equation \( HA\rightarrow\ H^+(aq)+A^-(aq)\) and create the appropriate balanced equation of your question. What is your acid or HA here? What is your conjugate base or A- here?

OpenStudy (andriani):



OpenStudy (andriani):

ummm... would u tell me the reason why we use this equipment??